The distance between Memphis and Atlanta is 782mi less than the distance between Minneapolis and Houston. Together, the two distances total 1570mi. How far is it between Memphis and Atlanta? How far is it between Minneapolis and Houston?

Answer by ReadingBoosters(3246)

x = Memphis to Atlanta
y = Minneapolis to Houston
x = y - 782
x + y = 1570
Substitute the first equation into the second equation,
y - 782 + y = 1570
Combine like terms
2y - 782 = 1570
2y - 782 + 782 = 1570 + 782
2y = 2352
2y
2y/2 = 2352/2
y = 1176
Plug y into the first equation and solve for x
x = 1176 - 782 = 394
Memphis and Atlanta 394 miles
Minneapolis and Houston 1176 miles